Hyper Drive Co is an early stage growth company and they need to plan for their capital requirements for the upcoming year. What working capital increases and PP&E investments should they be prepared for given the following given the revenue projections they have? Show work and please state any assumptions made. Current Year: Revenue: $350 million COGS: $195 million Working Capital: $80 million PP& E: $500 million Next Year: Sales Forecast Projects: $500 million in Revenue COGS will remain the same as a % of sales, as will Asset Turn.
Answer #1
Fig. in mlns. Current Yr Forecast Inc.%
Revenue 350 500 (500-350)/350= 42.86%
Less: COGS 195 278.57
Earnings 155 221.43
NWC 80 114.29
PP&E 500 714.29
580 828.57
ATO 0.6034 0.6034
Working capital & PPE increase with sales=Increase in sales %
ie. (500-350)/350= 42.86%
Increase in Working capital=80*0.4286= $ 34.29
Increase in PP&E =500*0.4286= $ 214.30